OpenStudy (anonymous):

(square root of 75b over c) plus (square root of 4b over 3c) plus (square root of 25b over c),

OpenStudy (anonymous):

i started this in previous post. but unclear what problem is. if you want look at previous one

OpenStudy (anonymous):

or not. i have to say it is somewhat annoying to start a problem and then see it posted again. but then again i am easily annoyed

OpenStudy (anonymous):

Surely not...:-)

OpenStudy (anonymous):

Why not take those common c's out from under the square root signs to start with?

OpenStudy (anonymous):

i wasn't sure if the "c" are underneath radical or not. if so then rationalize denominator and add. if not then just add.

OpenStudy (anonymous):

I can't get that link to work....

OpenStudy (anonymous):

OK, got it now...

OpenStudy (anonymous):

Ah, I see.

OpenStudy (anonymous):

u there, Paul?

OpenStudy (anonymous):

it is the previous post of this exact same problem. ok i must be getting grumpy so i will shut up and go do something else. but really, someone posts a problem, you spend some latex time asking for clarification, and then \[\color{red}{\text{BAM HERE IT IS AGAIN!}}\]

OpenStudy (anonymous):

Guess he's doing something else...

OpenStudy (anonymous):

satellite?

OpenStudy (anonymous):

U there now?

OpenStudy (anonymous):

i am here now.

OpenStudy (anonymous):

like telephone tag

OpenStudy (anonymous):

u know how java applets stop working when you leave the page for another tab, would be good if the people icon would disappear...

OpenStudy (anonymous):

ring ring. lets just do this \[\sqrt{\frac{75b}{c}}=\sqrt{\frac{75b}{c}}\times \frac{\sqrt{c}}{\sqrt{c}}=\frac{\sqrt{75bc}}{c}\]

OpenStudy (anonymous):

likewise \[\sqrt{\frac{4b}{3c}}=\frac{2\sqrt{3bc}}{3c}\]

OpenStudy (anonymous):

Guess he could manage the rest with that for an example.

OpenStudy (anonymous):

Oh, you're doing it all, sorry.

OpenStudy (anonymous):

and also \[\sqrt{\frac{25b}{c}}=\frac{5\sqrt{bc}}{c}\]

OpenStudy (anonymous):

tnx for the patience

OpenStudy (anonymous):

final work is to write \[\sqrt{75bc}=5\sqrt{3bc}\] get a least common denominator of c and 3c which is 3c, and add them up

OpenStudy (anonymous):

so get \[\frac{15\sqrt{3bc}+2\sqrt{3bc}+15\sqrt{bc}}{3c}\]

OpenStudy (anonymous):

unless i made a mistake. first two terms in the numerator are the same so you can add them. third is not

OpenStudy (anonymous):

okay thank you!